OpenStudy (gabylovesyou):

Identify which distributive property problem below would help you solve the following scenario: Your English class has 27 students. You would like to treat the class to cookies. You would like each person to have 3 cookies. Which expression would help you figure out how many cookies you would need to bake? 3(20 + 7) 3(27 + 3) 3(20 − 7) 3(27 − 3)

OpenStudy (theviper):

First

OpenStudy (theviper):

Because we want to treat everyone 3 cookies:- 3*27

OpenStudy (theviper):

& therefore first is right :P

OpenStudy (anonymous):

Wait, distributive property is helping in here. So 3(20+7) doesn't that make it 3(20) and 3(7)?

OpenStudy (gabylovesyou):

the choices are already in distributive

OpenStudy (gabylovesyou):

i guess lol

OpenStudy (theviper):

3(20+7) =3*27=81 by ur method 3(20)+3(7) 60+21 =81 :P

OpenStudy (anonymous):

Ah, right :P

OpenStudy (theviper):

27*3=81

OpenStudy (theviper):

\[\Huge{\color{red}{\text{:P}}}\]